OpenStudy (anonymous):

how to algebraically solve as lim=> infinity (3x+sin8x)/8x

OpenStudy (zzr0ck3r):

do you know lhopital rule?

OpenStudy (anonymous):

alas, no. my bc calc class hasn't gone over that yet.

OpenStudy (zzr0ck3r):

hmm nm this does not help...

OpenStudy (zzr0ck3r):

do you know about the squeeze theorem?

OpenStudy (anonymous):

yep! but I have to admit I have not used it a lot. I could understand maybe

OpenStudy (anonymous):

you can break it up... 3x/8x + (sin 8x)/8x first is 3/8, second is 0. yu can show this using the squeeze therom... -1 le sin 8x le 1 -1/8x goes to 0 as does 1/8x so (sin 8x)/8x must also go to 0.

OpenStudy (anonymous):

le is <=

OpenStudy (zzr0ck3r):

\[-1\le sin(8x)\le1 \implies-\frac{1}{x}\le \frac{sin(8x)}{x}\le\frac{1}{x}\implies \frac{sin(8x)}{x} \rightarrow 0\] so \[\frac{3x+sin(8x)}{3x}=\frac{\frac{3}{x}+\frac{sin(x)}{x}}{\frac{8x}{x}}=\frac{3+\frac{sin(8x)}{x}}{8}\rightarrow\frac{3}{8}\]

OpenStudy (anonymous):

thank you all.